14000 hired employees in 2015
2000 of the employees are being promoted in the year 2016 all receiving
$6000 sign on bonuses.
If there were 12500 newly hired employees in the year 2014 and of them only
1500 were promoted for the year 2015 still each receiving the $6000 bonuses
what would the bonus average be for each employee if the employer created a table for bonuses which is, "we will award bonuses on top of the sign on based on the percentage of the increase of the The amount we set aside for bonuses with be multiplied by the percentage increase of in house promotions this year as opposed to the year before and that percentage would equal what amount

Answer by JoelSchwartz(130)

x=the percent of increase for the total bonus in 2015 to 2016.
9,000,000x=12,000,000
x=12/9
y=the percent of increase for the amount of in house promotions.
1500y=2000
3y=4
y=4/3
(48/3)/9=16/9
Both percentages multiplied equal 16/9